OpenStudy (anonymous):

Mikalee, a fighter pilot, ejects from her burning plane. Once she is safely on the ground, she launches a flare to attract the attention of a rescue plane. The initial velocity of the flare is 1600 ft/sec. The height of the flare, h(t), at any time, t, is modeled by the function h(t) = - 16t^2 + 1600t. What is the maximum height of the flare? 50 ft 40000 ft 400 ft 4000 ft

OpenStudy (anonymous):

jeez h ow many more of these you got?

OpenStudy (anonymous):

20..

OpenStudy (anonymous):

max is at \(t=-\frac{b}{2a}=\frac{1600}{32}=50\) then compute \[f(50)=-16\times 50^2+1600\times 50=40,000\]

OpenStudy (anonymous):

So B?